# Text2Afford: Probing Object Affordance Prediction abilities of Language Models solely from Text

### v2

@@ Abstract
We investigate the knowledge of object affordances in pre-trained language models (LMs) and pre-trained Vision-Language models (VLMs). A growing body of literature shows that PTLMs fail inconsistently and non-intuitively, demonstrating a lack of reasoning and grounding. To take a first step toward quantifying the effect of grounding (or lack thereof), we curate a novel and comprehensive dataset of object affordances – Text2Afford, characterized by 15 affordance classes. Unlike affordance datasets collected in vision and language domains, we annotate in-the-wild sentences with objects and affordances. Experimental results reveal that PTLMs exhibit limited reasoning abilities when it comes to uncommon object affordances. We also observe that pre-trained VLMs do not necessarily capture object affordances effectively. Through few-shot fine-tuning, we demonstrate improvement in affordance knowledge in PTLMs and VLMs. Our research contributes a novel dataset for language grounding tasks, and presents insights into LM capabilities, advancing the understanding of object affordances.  [^1] Code and Data are available anonymously at https://anonymous.4open.science/r/Text2Afford
@@ 1 Introduction
Gaining the knowledge of object affordances while learning textual representation from large corpora maybe hard; as in NLP, we lack corresponding images (or videos) which provides necessary visual cues such as shape, color, and texture to predict affordances. This lack of mapping or rather grounding ability has been noted by many researchers in the context of large pretrained language models (PTLMs). Authors in Bender and Koller (2020) have pointed the lack of symbol grounding to be a fundamental factor behind PTLMs failing to grasp meaning from form (surface form text). The authors argue that language models which are exposed to only text (surface form) may never truly understand meaning, as PTLMs are unaware of possible groundings of the surface text.
@@
Table 1: Comparison of Text2Afford with other reasoning datasets. A larger version is in Appendix A.3 Table 9.
@@
$\bullet$ Using Text2Afford, we perform zero-shot evaluation of several state-of-the-art PTLMs along with a few VLMs in different settings to identify the extent to which they gain the knowledge of affordance during pretraining. We further ensemble the VLM and the PTLM predictions to examine whether pre-training with images can enrich affordance prediction from text. Overall, we observe that the SOTA LLMs face difficulties predicting contextual object affordances solely from text (accuracy $<55\%$) and the performance gets slightly enhanced when using powerful VLMs in presence of synthetic images.
@@ 2 Related work
Probing methods. Talmor et al. (2020) utilizes probing and employs Multi-choice MLM (Masked Language Modelling) and Multi-choice QA (Question Answering) setup to capture reasoning capabilities of pre-trained Language Models. Yang et al. (2022) examines zero-shot prediction performances on different tasks by LLM through novel visual imagination. Aroca-Ouellette et al. (2021) highlights the shortcomings of state-of-the-art pre-trained models in physical reasoning, with a further performance decline observed when incorporating option shuffling and superlatives in reasoning questions. Liu et al. (2022) proposes a novel spatial commonsense probing framework to investigate object scales and positional relationship knowledge in text-based pre-trained models and models with visual signals.
@@ 3 Text2Afford dataset construction
Text2Afford data exploration. We observe that classes such as ‘Grasp’, ‘Lift’, ‘Throw’, ‘Push’, and ‘Watch’ are the most common affordances for the objects present in the dataset (see Figure 2). Most frequent objects and their corresponding agreement scores are shown in Appendix A.10 Fig. 8. We observe, agreement scores are fairly uniform ($0.5$-$0.6$) for frequent objects, with high agreement for some frequent objects ($0.8$ for “the movie”). In Figure 9 (see Appendix A.9), we also see that ‘Grasp’, ‘Lift’, and ‘Throw’ are highly correlated classes. There is similar positive correlation between the class ‘SitOn’ and ‘Ride’, and some correlation between ‘Watch’ and ‘LookThrough’.
@@ 6 Benchmarking Text2Afford prediction
FLAN-T5-large 0.06 0.13±$\pm 0.04$
FLAN-T5-xl 0.07 0.21±$\pm 0.03$
FLAN-T5-xxl 0.33 0.39±$\pm 0.04$
Falcon-7b-instruct 0.19 0.24±$\pm 0.03$
Falcon-40b-instruct 0.43 0.47±$\pm 0.06$
Llama-3-8b-instruct 0.36 0.43±$\pm 0.05$
ChatGPT (GPT-3.5 turbo) 0.41 0.44±$\pm 0.05$
@@
Idefics-9b-instruct 0.26 0.25 0.36±$\pm 0.02$ 0.37±$\pm 0.03$
llava-1.5-7b 0.32 0.34 0.36±$\pm 0.03$ $\pm 0.04$
instructblip-vicuna-13b 0.37 0.39 0.43±$\pm 0.03$ 0.45±$\pm 0.03$
instructblip-flan-t5-xl 0.12 0.16 0.15±$\pm 0.02$ 0.18±$\pm 0.02$
instructblip-flan-t5-xxl 0.39 0.45 0.48±$\pm 0.04$ 0.53±$\pm 0.05$
@@
{mdframed}
[backgroundcolor=gray!10] Sentence: The salt from La Mata is often used as table salt. Object: table salt
Top 5 predicted affordances (according to the probability score) - [‘sitOn’, ‘pourFrom’, ‘grasp’, ‘fix’, ‘lookThrough’]
@@ 7 Text2Afford for physical reasoning
FLAN-T5-xxl 0.34 0.38(+.04) 0.72 0.75(+.03)
Table 8: Text-only physical reasoning dataset evaluation using different LLMs fine-tuned on Text2Afford. +Text2Afford: instruction fine-tuned on Text2Afford
@@ A.1 Details of the Text2Afford dataset construction
Pilot studies & annotator training. We annotate the dataset using the Toloka platform [^9] https://toloka.ai/. We design an interface on this platform, which contained clear instructions and examples for annotating the data. We conduct two rounds of pilot studies to analyze the subjective understanding of the annotators and, thereby, filter out the high quality, serious annotators. For the first pilot study, we present the annotators with the smaller 62 sentence-object pairs and ask them to label the instance with each affordance class on a scale of 00 to $3$, indicating whether or not the affordance can be performed on the object. Here, 0-1 indicates that the affordance cannot be performed (high-low) and 2-3 indicates that the affordance can be performed low-high). We will further use these 62 synthetic sentence-object pairs for few-shot training. For quality control, we select the top 90% of the available annotators in the platform, who are proficient in English, and use computers to complete the tasks [^10] We exclude mobile-users as we believe the instructions may not appear clearly on mobile devices.. A total of 15 annotators labelled the data, and all of them were incentivized uniformly. After the first pilot, we find that there is an extremely poor agreement among the annotators, and the overall precision is around 28%. Therefore, we moved on to a second pilot study. Here, we use all the 62 sentence-object pairs from the previous study, along with 32 randomly selected sentence-object pairs from the XNLI data. We use the top 30% of the annotators (based on the quality determined by the platform) available on the platform, while other criteria remained the same. We annotate 32 sentence-object pairs ourselves, and use all the labelled examples as control data points to guide the annotators while labelling. A total of 114 annotators (including the 14 annotators from the first pilot study) participated in this version of the pilot study. We assign a specific skill to the annotators who attained more than $30\%$ precision and $30\%$ recall. In total, 48 annotators passed this criteria. Through initial pilot studies, we learnt that without grounded images, the task appears quite subjective to annotators. The main goal of the pilot studies have been to understand the annotators’ quality, their comprehension of the task, and their preferences for incentives per task. We have also conducted two additional AMA (Ask Me Anything) sessions with interested annotators to further clarify the task.
@@ A.2 Additional analysis on the datapoints by human
To further interpret the difficulty (or ambiguity) of the datapoints, we filter out the “sentence-object-affordance” triples based on the percentage annotator agreement. We categorize the triples into 3 sections: {mdframed} Agreement > 0.75: Total 26,411 triples 0.4<Agreement<0.75: Total 7,084 triples Agreement < 0.4: Total 2,025 triples In general, the average agreement is higher for negative affordance classes than that of positive classes, which implies that it is easier for humans to tell which ‘affordance’ is not applicable to a particular object.
@@
A.3 Comparison with other reasoning dataset
$\alpha$NLI Bhagavatula et al. (2020) 169,654 - 1,532 Abductive logical reasoning Crowd-sourced
$\alpha$ARCT Niven and Kao (2019) 2420 632 888 Abductive logical reasoning Crowd-sourced
Table 9: Comparison of Text2Afford with other reasoning datasets.
A.4 Instruction page on the Toloka platform
@@ A.6 Annotators demographics
A.7 Phasewise annotator agreement
@@ A.9 Correlation of affordances
A.10 Most frequent objects
@@ Appendix B Experimental setup
B.1 Zero-shot experiments
B.1.1 Commonsense reasoning tasks
@@
B.2 Ensemble language and vision prediction
@@ B.3 Few-shot experiments
B.3.1 Multimodal task setup
@@ Appendix D Prompt selection
Appendix E Model implementation details
@@
Appendix F Details of evaluation metric
@@
Appendix G Dataset creation time
@@
Appendix H Sample dataset
@@
Appendix I Additional experiments
I.1 Qualitative analysis of generated images

### v3

@@ Abstract
We investigate the knowledge of object affordances in pre-trained language models (LMs) and pre-trained Vision-Language models (VLMs). A growing body of literature shows that PTLMs fail inconsistently and non-intuitively, demonstrating a lack of reasoning and grounding. To take a first step toward quantifying the effect of grounding (or lack thereof), we curate a novel and comprehensive dataset of object affordances – Text2Afford, characterized by 15 affordance classes. Unlike affordance datasets collected in vision and language domains, we annotate in-the-wild sentences with objects and affordances. Experimental results reveal that PTLMs exhibit limited reasoning abilities when it comes to uncommon object affordances. We also observe that pre-trained VLMs do not necessarily capture object affordances effectively. Through few-shot fine-tuning, we demonstrate improvement in affordance knowledge in PTLMs and VLMs. Our research contributes a novel dataset for language grounding tasks, and presents insights into LM capabilities, advancing the understanding of object affordances.  [^1] Code and Data are available at https://github.com/sayantan11995/Text2Afford
@@ 1 Introduction
Object affordance refers to the properties of an object that determine what actions a human can perform on them Gibson (1979). Gaining the knowledge of object affordances while learning textual representation from large corpora maybe hard; as in NLP, we lack corresponding images (or videos) which provides necessary visual cues such as shape, color, and texture to predict affordances. This lack of mapping or rather grounding ability has been noted by many researchers in the context of large pretrained language models (PTLMs). Authors in Bender and Koller (2020) have pointed the lack of symbol grounding to be a fundamental factor behind PTLMs failing to grasp meaning from form (surface form text). The authors argue that language models which are exposed to only text (surface form) may never truly understand meaning, as PTLMs are unaware of possible groundings of the surface text.
@@
Table 1: Comparison of Text2Afford with other reasoning datasets.
@@
$\bullet$ Using Text2Afford, we perform zero-shot evaluation of several state-of-the-art PTLMs along with a few VLMs in different settings to identify the extent to which they gain the knowledge of affordance during pretraining. We further ensemble the VLM and the PTLM predictions to examine whether pre-training with images can enrich affordance prediction from text. Overall, we observe that the SOTA LLMs face difficulties predicting contextual object affordances solely from text (accuracy $55\%$) and the performance gets slightly enhanced when using powerful VLMs in presence of synthetic images.
@@ 2 Related work
Probing methods. Talmor et al. (2020) utilizes probing and employs Multi-choice MLM (Masked Language Modelling) and Multi-choice QA (Question Answering) setup to capture reasoning capabilities of pre-trained Language Models. Yang et al. (2022) examines zero-shot prediction performances on different tasks by LLM through novel visual imagination. Aroca-Ouellette et al. (2021) highlights the shortcomings of state-of-the-art pre-trained models in physical reasoning, with a further performance decline observed when incorporating option shuffling and superlatives in reasoning questions. Liu et al. (2022) proposes a novel spatial commonsense probing framework to investigate object scales and positional relationship knowledge in text-based pre-trained models and models with visual signals. Joshi et al. (2020) uses probing methods to investigate a more fine-grained logical reasoning capabilities of pre-trained models.
@@ 3 Text2Afford dataset construction
Text2Afford data exploration. We observe that classes such as ‘Grasp’, ‘Lift’, ‘Throw’, ‘Push’, and ‘Watch’ are the most common affordances for the objects present in the dataset (see Figure 2). Most frequent objects and their corresponding agreement scores are shown in Appendix A.9 Fig. 8. We observe, agreement scores are fairly uniform (0.50.5-$0.6$) for frequent objects, with high agreement for some frequent objects ($0.8$ for “the movie”). In Figure 9 (see Appendix A.8), we also see that ‘Grasp’, ‘Lift’, and ‘Throw’ are highly correlated classes. There is similar positive correlation between the class ‘SitOn’ and ‘Ride’, and some correlation between ‘Watch’ and ‘LookThrough’.
@@ 6 Benchmarking Text2Afford prediction
Random baseline 0.18 -
[…]
@@
[…]
@@
Sentence: The salt from La Mata is often used as table salt. Object: table salt Top 5 predicted affordances (according to the probability score) - [‘sitOn’, ‘pourFrom’, ‘grasp’, ‘fix’, ‘lookThrough’]
@@ 7 Text2Afford for physical reasoning
FLAN-T5-xxl 0.34 0.38(+.04)* 0.72 0.75(+.03)
Table 8: Text-only physical reasoning dataset evaluation using different LLMs fine-tuned on Text2Afford. +Text2Afford: instruction fine-tuned on Text2Afford. * indicates $p$-value (<0.05<0.05) using Mann-Whitney U-Test.
8 Additional details
Reason for choosing XNLI.
We select XNLI to incorporate object references from less conventional and commonly explored scenarios. Unlike typical object identification datasets, XNLI offers sentences derived from novels, thus presenting a more in-the-wild textual context, which adds complexity and diversity to our dataset. Specifically, we choose the hypothesis portion of the XNLI sentences due to its shorter context length. This choice intentionally poses a challenge to LLMs, allowing us to better evaluate their reasoning capabilities, especially when dealing with minimal contextual information.
Non-explicit mention about contextual object affordance in the instruction.
The instructions shown in Appendix Figure 3 represent the initial guidance provided to annotators as an introduction to the task. Since understanding contextual object affordance can be challenging for non-expert annotators, this initial step was designed to give a basic idea of the task. However, we follow this up with a comprehensive training process and conduct two AMA (Ask Me Anything) sessions to ensure that annotators fully understood the need to base their judgments on the provided context. These efforts are key in ensuring high-quality annotations throughout the dataset creation.
Reason for choosing 0-3 Likert scale in data annotation.
We opt for a 0-3 Likert scale (4-point) to minimize the potential for neutral or non-committal responses, which can often arise when a midpoint option is available. Our initial observations indicated that some annotators tended to select an “average" value without fully considering the contextual affordance of the objects, which diminished the depth of their evaluations and limited the discussion around ambiguities. By adopting a 4-point scale, we aim to encourage more decisive judgments. In addition, we provide a textbox (see Appendix Figure 3) for annotators to express any uncertainties or ambiguities they encountered, which has helped us in capturing more nuanced feedback.
Reason for choosing visual genome.
We chose visual genome as a primary source for real images due to its rich, complex scenes, which are widely used in visual reasoning tasks. The complexity of the images in visual genome provides diverse contexts that align well with the goals of our study, which focuses on contextual object affordances. While other methods, such as using search engines like Bing, have been employed in prior work to retrieve images, we opt for visual genome to ensure that the images contain sufficient contextual and visual detail to support affordance prediction, even if there are minor limitations in reasoning.
Reason for choosing stable diffusion.
Regarding the use of stable diffusion, we have been inspired by its demonstrated capability to generate high-quality, realistic images, particularly in prior studies where it was effective in reasoning tasks. While CLIP is primarily trained on real-world images, we hypothesize that stable diffusion could generate contextual images with sufficient accuracy to complement the real images. The generated images provide additional diversity, which helps us explore the affordance prediction task from a different angle. The benefit of using stable diffusion lies in its ability to create controlled, context-specific images that may not always be available in existing datasets, providing a broader range of testing scenarios for our models.
Reason for framing generative tasks as a binary decision problem.
In the generative setting, we opt for a binary yes/no classification to evaluate the affordance of individual context-object-affordance triples. We decide this based on the observation of the tendency of smaller LLMs to hallucinate, which can make direct affordance prediction challenging, particularly in zero-shot scenarios. By framing it as a binary classification task, we aim to simplify the evaluation and obtain more reliable results. In addition, our approach allows for a comprehensive evaluation of both positive and negative affordances. This is critical for our dataset, as it is designed to assess affordances that are applicable, as well as those that are not, in a given context.
@@ 9 Conclusion
Acknowledgments
We would like to express our sincere gratitude to our co-authors for their invaluable contributions throughout this work. We also extend our thanks to the reviewers for their constructive feedback, which significantly helped improve the quality of the paper. Additionally, we gratefully acknowledge the support of the Toloka Research Grant program, which partially funded the data annotation process.
@@ A.1 Details of the Text2Afford dataset construction
Pilot studies & annotator training. We annotate the dataset using the Toloka platform [^9] https://toloka.ai/. We design an interface on this platform, which contained clear instructions and examples for annotating the data. We conduct two rounds of pilot studies to analyze the subjective understanding of the annotators and, thereby, filter out the high quality, serious annotators. For the first pilot study, we present the annotators with the smaller 62 sentence-object pairs and ask them to label the instance with each affordance class on a scale of 0 to $3$, indicating whether or not the affordance can be performed on the object. Here, 0-1 indicates that the affordance cannot be performed (high-low) and 2-3 indicates that the affordance can be performed low-high). We will further use these 62 synthetic sentence-object pairs for few-shot training. For quality control, we select the top 90% of the available annotators in the platform, who are proficient in English, and use computers to complete the tasks [^10] We exclude mobile-users as we believe the instructions may not appear clearly on mobile devices.. A total of 15 annotators labelled the data, and all of them were incentivized uniformly. After the first pilot, we find that there is an extremely poor agreement among the annotators, and the overall precision is around 28%. Therefore, we moved on to a second pilot study. Here, we use all the 62 sentence-object pairs from the previous study, along with 32 randomly selected sentence-object pairs from the XNLI data. We use the top 30% of the annotators (based on the quality determined by the platform) available on the platform, while other criteria remained the same. We annotate 32 sentence-object pairs ourselves, and use all the labelled examples as control data points to guide the annotators while labelling. A total of 114 annotators (including the 14 annotators from the first pilot study) participated in this version of the pilot study. We assign a specific skill to the annotators who attained more than $30\%$ precision and $30\%$ recall. In total, 48 annotators passed this criteria. Through initial pilot studies, we learnt that without grounded images, the task appears quite subjective to annotators. The main goal of the pilot studies have been to understand the annotators’ quality, their comprehension of the task, and their preferences for incentives per task. We have also conducted two additional AMA (Ask Me Anything) sessions with interested annotators to further clarify the task.
@@ A.2 Additional analysis on the datapoints by human
To further interpret the difficulty (or ambiguity) of the datapoints, we filter out the “sentence-object-affordance” triples based on the percentage annotator agreement. We categorize the triples into 3 sections: Agreement >0.75>0.75: Total 26,411 triples 0.4<Agreement <0.75<0.75: Total 7,084 triples Agreement <0.4<0.4: Total 2,025 triples In general, the average agreement is higher for negative affordance classes than that of positive classes, which implies that it is easier for humans to tell which ‘affordance’ is not applicable to a particular object.
@@
A.3 Instruction page on the Toloka platform
@@ A.5 Annotators demographics
A.6 Phasewise annotator agreement
@@ A.8 Correlation of affordances
A.9 Most frequent objects
@@ Appendix B Experimental setup
B.1 Random baseline
In addition to evaluating generative models, we establish a random baseline. For this baseline, we randomly assign "yes" to the 15 affordance classes for each sentence-object pair, with random selections made from 0 to 9 (based on the observation that the maximum number of positive affordances per pair is 9). Interestingly, we find that models like Flan-T5-large and Flan-T5-XL underperform compared to this random baseline in zero-shot settings, highlighting the inherent difficulty of the task in such scenarios.
B.2 Zero-shot experiments
B.2.1 Commonsense reasoning tasks
@@
B.3 Ensemble language and vision prediction
@@ B.4 Few-shot experiments
B.4.1 Multimodal task setup
@@ Appendix D Prompt selection
Appendix E Instruction fine-tuning setup
Data sample selection.
We select sentence-object pairs from the Text2Afford dataset where at least one positive affordance is present. For each selected sentence-object pair, we randomly assign one positive affordance and one negative affordance, yielding a balanced dataset of 1819 training instances (positive and negative classes). To incorporate additional domain knowledge and reduce the likelihood of generating hallucinated answers, we include 500 randomly sampled instances from the training set of the target task (i.e., PIQA). For the PROST task, as the training set is not explicitly available, we sample from the test set and ensure these samples are removed from the evaluation set during testing. The training instances are framed in a multiple-choice question answering format.
Fine-tuning setup.
We utilize Alpaca-formatted prompts (shown in Table 10, Table 11 and Table 12 for the Text2Afford, PIQA and PROST tasks, respectively). We fine-tune 4-bit quantized models with PEFT, focusing on the adapter layers. We perform the fine-tuning over 5 epochs with a batch size of 8, a learning rate of 2e-10, weight decay, and a maximum sequence length of 256.
Instruction to fine-tune Text2Afford Below is an instruction that describes a task. Write a response that appropriately completes the request. ### Instruction:
You are an AI assistant that has strong reasoning capability. You are given a context containing an object, and you are asked to answer a question about the object based on the context. Just response ’Yes’ or ’No’.
### Context:
{context}
### Object:
{object}
### Question:
Can human {affordance} the {object}?
### Answer:
{answer}
Table 10: Instruction to fine-tune Text2Afford.
Instruction to fine-tune PIQA Below is an instruction that describes a task. Write a response that appropriately completes the request. ### Instruction:
You are an AI assistant that has strong reasoning capability. You are given a situation and asked to choose the most appropriate option from given two options.
### Situation:
{situation}
### Options:
[0]
{option0}
[1] {option1}
Only response the ‘answer id’. For example if the answer is [0] then response 0. DO NOT respond anything other than <0, 1>. ### Answer:
{answer}
Table 11: Instruction to fine-tune PIQA.
Instruction to fine-tune PROST Below is an instruction that describes a task. Write a response that appropriately completes the request. ### Instruction:
You are an AI assistant that has strong reasoning capability. You are given a question with 4 options and you have to choose the right option.
### Question:
{question}
### Options:
[0]
{option_A}
[1] {option_B}
[2] {option_C}
[3] {option_D}
Only response the ‘answer id’. For example if the answer is [0] then response 0. DO NOT respond anything other than <0, 1, 2, 3>. ### Answer:
{answer}
Table 12: Instruction to fine-tune PROST.
Appendix F Model implementation details
@@
Appendix G Details of evaluation metric
@@
Appendix H Dataset creation time
@@
Appendix I Sample dataset
@@
Appendix J Additional experiments
J.1 Qualitative analysis of generated images